The Length of a Rectangular garden is 2 more than twice the width, and the area is 40ft^2. What are the dimensions W and L?

Answer by ewatrrr(24785)

 
Hi,

Let x represent the width
Question states***
(2x+2)x = 40ft^2
2x^2 + 2x - 40 = 0
x^2 + x - 20 = 0
(x+5)(x-4) = 0 (Tossing out the negative solution for unit measure)
x = 4ft, the width. The length is 10ft